OpenStudy (anonymous):

What is the solution to the equation? t(9 – 1) = 5 • 9 – 5 • 1 A. –1 B. 1 C. 9 D. 5

OpenStudy (kirbykirby):

t(9-1)=(5)(9)-(5)(1) t(8)=45-5 8t=40 t=40/8 t=5

OpenStudy (anonymous):

Ooooooooh. I was thinking that it was the answer to equation.

OpenStudy (anonymous):

Thank you for helping me understand it.

OpenStudy (kirbykirby):

:)

OpenStudy (anonymous):

I have another question. Can you help me with it?

OpenStudy (kirbykirby):

Sure?